# Hongyi Du1, Jiarui Yao1, Xiaocheng Yang1, Denghui Zhang1,3, Yunzhu Li4, Heng Ji1 1

### v1

@@ 5.1 Settings
Table 3: Benchmarking results of baseline BaseAgent with different core models on EscapeBench. Both closed- and open-source models rely heavily on hints to complete the escape compared to human performance, with smaller-scale models exhibiting a particularly high dependency.
@@ 5.2 Benchmarking Results
We benchmark current models using the BaseAgent framework, with results in Table 3 showing that large-scale closed-source models, such as GPT-4o, Claude-3.5, and Gemini-1.5, consistently outperform smaller models. Key insights include:
@@
  • Most hints are used on key steps, which require creative reasoning, while models can often collect tools through random exploration.
@@
  • Models require significantly more action steps and hints than the average human player, with some using up to  20x more steps than the most efficient action chain.
Figure 5: Ablation of difficulty through Hints Used. Harder game settings generally require more hints to complete, and vice versa.
Figure 6: Ablation of difficulty through Total Steps. Harder game settings generally require more total steps to complete, and vice versa.
Complex environment causes challenges. We show in Figure 5 and Figure 6 that, generally, harder game settings require more steps and hints for an agent to solve. Since our difficulty setting depends solely on the granularity and usefulness of descriptions and feedback (see Table 1), our results demonstrate that the way the environment is presented can impact difficulty, even when the core game logic remains unchanged.
Figure 7: Distribution of Key Steps Hints Used, categorized by different actions. Colored bars represent the percentage of hints used for each action type relative to the total key steps for that type (See right of Figure 3). “Craft” and “Input” account for the highest percentage of relative hint usage, highlighting their challenge.
@@
Input and craft are the most challenging. As shown in Figure 7, while “Apply” actions require the most hints overall, “Input” and “Craft” actions show the highest relative hint usage compared to their total occurrence in key steps. This reflects the “Input” action’s large parametric space, making random guesses impractical. Additionally, the implicit numerical challenges presented by “Input” often require synthesizing multiple pieces of evidence, adding to their difficulty. The results also suggest that the model demonstrates a degree of creativity, enabling effective “Apply” of tools, but “Craft” actions are inherently more creativity-demanding.
Table 4: Error Analysis of BaseAgent’s inefficiency or failures. All cases are selected in the same game setting. Superficial attempts dominate larger-scale models, inspiring EscapeAgent’s design to address them.
@@
Error analysis. We observe from Table 4 that the BaseAgent often gets stuck and relies on hints due to its struggles with environment-following and creativity. Smaller models tend to perform invalid actions in complex scenarios, while larger models excel at tool collection but fail to attempt creative actions, resorting to superficial strategies. Our EscapeAgent addresses this by promoting more purposeful and creative actions.
GPT-4o $5.03_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.27}}$ $452.75_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 270.86}}$ $47.03_{{\color[rgb]{1,0,0}\uparrow 22.28}}$ $0.33_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 1.84}}$ (1.19%) $4.70_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 3.44}}$ (13.74%)
GPT-4o-mini $10.58_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 4.61}}$ $752.25_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 250.14}}$ $28.17_{{\color[rgb]{1,0,0}\uparrow 12.11}}$ $1.14_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 0.86}}$ (4.16%) $9.44_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 3.75}}$ (28.53%)
Llama-3.1-70B-Instruct $7.92_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 6.61}}$ $645.19_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 337.23}}$ $31.44_{{\color[rgb]{1,0,0}\uparrow 12.44}}$ $1.42_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 1.69}}$ (5.44%) $6.56_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 4.86}}$ (19.15%)
Qwen2.5-72B-Instruct $9.72_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 6.78}}$ $746.61_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 355.89}}$ $28.62_{{\color[rgb]{1,0,0}\uparrow 16.16}}$ $2.72_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 2.61}}$ (10.61%) $7.00_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 4.17}}$ (20.71%)
DeepSeek-LLM-67b-Chat $20.14_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.36}}$ $1285.03_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 273.44}}$ $15.30_{{\color[rgb]{1,0,0}\uparrow 8.67}}$ $10.81_{{\color[rgb]{1,0,0}\uparrow 0.31}}$ (42.72%) $9.34_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.66}}$ (27.20%)
Yi-1.5-34B-Chat $22.59_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 1.41}}$ $1468.03_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 105.30}}$ $12.04_{{\color[rgb]{1,0,0}\uparrow 0.08}}$ $10.42_{{\color[rgb]{1,0,0}\uparrow 2.31}}$ (41.61%) $12.19_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 3.73}}$ (35.71%)
Phi-3-medium-128k-instruct $25.75_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 6.44}}$ $1513.69_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 357.50}}$ $9.99_{{\color[rgb]{1,0,0}\uparrow 2.65}}$ $11.20_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 0.91}}$ (44.99%) $14.55_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.56}}$ (43.35%)
Llama-3.1-8B-Instruct $19.81_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 6.05}}$ $1271.53_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 271.77}}$ $19.22_{{\color[rgb]{1,0,0}\uparrow 9.12}}$ $6.64_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 0.17}}$ (25.23%) $13.22_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.89}}$ (39.34%)
Ministral-8B-Instruct $19.47_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.84}}$ $1233.72_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 323.25}}$ $19.34_{{\color[rgb]{1,0,0}\uparrow 10.37}}$ $6.61_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 0.56}}$ (25.56%) $12.86_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 5.33}}$ (38.93%)
Qwen2.5-7B-Instruct $27.53_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 4.67}}$ $1639.58_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 310.84}}$ $8.56_{{\color[rgb]{1,0,0}\uparrow 2.04}}$ $13.00_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 0.81}}$ (52.66%) $14.58_{{\color[rgb]{0.0390625,0.640625,0.265625}\downarrow 3.89}}$ (43.99%)
Table 5: Benchmarking Results of our approach EscapeAgent with different core models. Nearly all the performance raises compared to Table 3, showcasing the effectiveness of EscapAgent in promoting the agent’s creativity.
@@ 5.3 EscapeAgent Results
The introduction of the Reflection and Foresight modules in EscapeAgent, as shown in Table 5, leads to a marked reduction in hint usage and total steps, with large-scale models benefiting the most. Key insights include:
  • Larger models like GPT-4o consistently outperform smaller models, suggesting that while the new modules aid creative reasoning, the core model’s capabilities remain a crucial factor.
  • Early exit progress improves significantly across models, despite the exponentially increasing difficulty of maintaining consecutive meaningful progress without hints. Notably, GPT-4o achieves near-human-level performance, demonstrating the effectiveness of the enhanced framework.
@@
EscapeAgent progresses more efficiently. Figure 9 demonstrates EscapeAgent’s more efficient progress, as the slopes of the dotted lines are consistently steeper. We further show in Figure 8 that EscapeAgent requires fewer actions to achieve the next valid Key Steps compared to BaseAgent, indicating better creative reasoning. The spike at 50 steps corresponds to explicit hints provided after the model’s prolonged inactivity for consecutively 50 steps, and the lower pink bar here further highlights EscapeAgent’s reduced hint dependency. Interestingly, for all models, sudden progress after 15 steps without hints is rare, highlighting a lack of spontaneous insights that humans often demonstrate, even in the EscapeAgent framework.
Figure 10: A comparison of action trial times distribution for a specific item before success. EscapeAgent tries more actions with a higher overall success rate.
EscapeAgent tries more with higher overall success. In Figure 10, we analyze the valid actions agents attempt on different items in scenes before successfully operating them. While the BaseAgent exhibits a higher first-attempt success rate, EscapeAgent achieves greater effectiveness by making multiple attempts, leading to a significantly higher overall success rate within 10 trials. This difference can be attributed to EscapeAgent’s strategy of proposing multiple viable actions simultaneously. Although this increases the likelihood of eventually succeeding by trying more actions, it does not prioritize the most-likely-to-succeed action first. As a result, the BaseAgent appears more efficient on its initial trial, but its superficial approach, as highlighted in Table 4, limits its overall performance. The EscapeAgent’s design effectively addresses this limitation by leveraging a more exploratory approach, which proves advantageous in complex scenarios requiring creativity.
@@
Case study. Figure 11 presents a case study of agent progress relative to action steps across six game settings. Red dots indicate steps where hints were provided. We observe that i) EscapeAgent requires significantly fewer hints and achieves a steeper progress curve; ii) EscapeAgent demonstrates the ability to make consecutive progress in shorter intervals (e.g., the light blue and pink lines for both GPT-4o and Llama-3.1-70B); iii) harder scenarios remain challenging, especially for BaseAgent, which often relies heavily on hints to make progress (e.g., the later part of the yellow line for GPT-4o); and iv) Average human still performs far better than all agents. Humans rarely make mistakes shown in Table 4, while both agents still struggle with short memory due to context length and creative tool use strategies. These findings emphasize the need for further improvements and highlight EscapeBench’s challenge to even the most advanced models.

### v2

@@ 4.3 Foresight Module
Table 3: Benchmarking results of BaseAgent with different core models on EscapeBench. An oracle action chain’s total step is only 107.83 on average. Both closed- and open-source models rely heavily on hints to complete the escape compared to human performance, with smaller-scale models exhibiting a particularly high dependency.
5 Experiments
We divide experiments into: i) Benchmarking model creativity within the BaseAgent, and ii) Evaluating EscapeAgent’s effectiveness.
5.1 Settings
Environment.
Experiments are conducted on 36 game settings. An agent is considered to be making progress if it either achieves a key step (defined in Figure 3) or collects a new tool. Agents will receive help if they fail to make progress for 50 consecutive actions (see Section D.1), thus ensuring full completion of the game. The working memory length is set to 10.
Models.
We evaluate both closed- and open-source models: GPT-4o, GPT-4o-mini (Hurst et al., 2024), Claude-3.5 (Anthropic, 2024), Gemini-1.5 (Team et al., 2024), Llama-3.1 (Dubey et al., 2024), Qwen-2.5 (Team, 2024), DeepSeek-LLM (Liu et al., 2024a), Phi-3.5 (Abdin et al., 2024), Yi (Young et al., 2024), Ministral (MistralAI, 2024). Models with fewer than 7B parameters are excluded due to near-random behavior. For consistency, we set sampling temperature to $T=0$ and $n=1$.
Metrics.
We use two main metrics including:
  • Hints Used: Total hints used in a game.
  • Total Steps: Total actions taken in a game.
Auxiliary metrics for analysis include:
Results are micro-averaged across the 36 settings.
Figure 5: Distribution of Key Steps Hints Used, categorized by different actions. Colored bars represent the percentage of hints used for each action type relative to the total key steps for that type (See right of Figure 3).
Table 4: Error Analysis of BaseAgent’s inefficiency or failures. All cases are selected in the same game setting.
GPT-4o $5.03_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 5.27}}$ $452.75_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 270.86}}$ $47.03_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 22.28}}$ $0.33_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 1.84}}$ (1.19%) $4.70_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 3.44}}$ (13.74%)
GPT-4o-mini $10.58_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 4.61}}$ $752.25_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 250.14}}$ $28.17_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 12.11}}$ $1.14_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 0.86}}$ (4.16%) $9.44_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 3.75}}$ (28.53%)
Llama-3.1-70B-Instruct $7.92_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 6.61}}$ $645.19_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 337.23}}$ $31.44_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 12.44}}$ $1.42_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 1.69}}$ (5.44%) $6.56_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 4.86}}$ (19.15%)
Qwen2.5-72B-Instruct $9.72_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 6.78}}$ $746.61_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 355.89}}$ $28.62_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 16.16}}$ $2.72_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 2.61}}$ (10.61%) $7.00_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 4.17}}$ (20.71%)
DeepSeek-LLM-67b-Chat $20.14_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 5.36}}$ $1285.03_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 273.44}}$ $15.30_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 8.67}}$ $10.81_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 0.31}}$ (42.72%) $9.34_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 5.66}}$ (27.20%)
Yi-1.5-34B-Chat $22.59_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 1.41}}$ $1468.03_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 105.30}}$ $12.04_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 0.08}}$ $10.42_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 2.31}}$ (41.61%) $12.19_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 3.73}}$ (35.71%)
Phi-3-medium-128k-instruct $25.75_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 6.44}}$ $1513.69_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 357.50}}$ $9.99_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 2.65}}$ $11.20_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 0.91}}$ (44.99%) $14.55_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 5.56}}$ (43.35%)
Llama-3.1-8B-Instruct $19.81_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 6.05}}$ $1271.53_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 271.77}}$ $19.22_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 9.12}}$ $6.64_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 0.17}}$ (25.23%) $13.22_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 5.89}}$ (39.34%)
Ministral-8B-Instruct $19.47_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 5.84}}$ $1233.72_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 323.25}}$ $19.34_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 10.37}}$ $6.61_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 0.56}}$ (25.56%) $12.86_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 5.33}}$ (38.93%)
Qwen2.5-7B-Instruct $27.53_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 4.67}}$ $1639.58_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 310.84}}$ $8.56_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\uparrow 2.04}}$ $13.00_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 0.81}}$ (52.66%) $14.58_{{\color[rgb]{0.0390625,0.640625,0.265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0390625,0.640625,0.265625}\downarrow 3.89}}$ (43.99%)
Table 5: Benchmarking Results of EscapeAgent with different core models. Nearly all the performance raises compared to Table 3, showcasing the effectiveness of EscapAgent in promoting the agent’s creativity.
@@ 5.2 Benchmarking Results
We benchmark current models using the BaseAgent framework, with results in Table 3 showing that large-scale closed-source models consistently outperform smaller models. Key insights include:
@@
  • Most hints are used on key steps, which demand creative reasoning, while models may often collect tools through random exploration.
@@
  • Models require significantly more action steps and hints than the average human, and up to 2̃0x more steps than the most efficient action chain.
We further present an ablation study of total hints and steps used in Section D.3.
Input and craft are the most challenging actions.
@@
As shown in Figure 5, while “Apply” actions require the most hints in absolute terms, “Input” and “Craft” actions have the highest relative hint usage compared to the total number of key steps. This likely reflects the large parametric space of “Input” actions, where random guesses are impractical, and the creativity-demanding nature of “Craft” actions.
Error analysis.
@@
We observe from Table 4 that the BaseAgent often gets stuck and relies on hints due to its struggles with environment-following and creativity. Smaller models tend to perform invalid actions in complex scenarios, while larger models excel at tool collection but fail to attempt creative actions, resorting to superficial strategies. Our EscapeAgent addresses this by promoting more purposeful and creative actions.
@@ 5.3 EscapeAgent Results
The introduction of the Reflection and Foresight in EscapeAgent, as shown in Table 5, significantly reduces hint uses and total steps, with larger models benefiting the most. Key insights include:
  • Larger models still outperform smaller ones, suggesting while new modules aid creative reasoning, the core model’s capabilities remain crucial.
  • Early exit progress improves across models, despite the exponentially increasing difficulty of consecutively making progress without hints, demonstrating EscapeAgent’s effectiveness.
@@
As shown in Figure 6, EscapeAgent demonstrates steeper progress slopes, reflecting greater efficiency. Figure 8 further shows that EscapeAgent requires fewer actions to reach the next key step, indicating stronger creative reasoning ability. The spike at 50 steps corresponds to hints provided after prolonged inactivity, so the lower pink bar here further highlights EscapeAgent’s reduced hint dependency. Notably, across all models, progress after 15 steps without hints is rare, underscoring a lack of spontaneous insights typically seen in humans.
Case study.
@@
Figure 7 illustrates agent progress relative to action steps across six game settings, with red dots marking hint provided. We observe that: i) EscapeAgent requires fewer hints and achieves steeper progress; ii) it can make consecutive progress in shorter intervals; iii) harder scenarios remain challenging, especially for BaseAgent, which heavily relies on hints; iv) Average human performs far better. Humans rarely make mistakes shown in Table 4, while agents still struggle with short memory due to context length and creative tool use strategies. These emphasize the need for further improvements and highlight EscapeBench’s challenge to even the most advanced models.
6 Further Analysis
6.1 Evaluation across Diverse Models
To enhance the diversity of model scales and domains in our evaluation, we present additional experimental results on smaller-scale and domain-specific models. These results supplement our main findings and provide further insights into the scalability and generalizability of our framework.
In addition to the models presented in out main table, we also evaluated sub-7B models, including the latest Llama-3.2-1B and 3B variants. However, these models demonstrated extremely limited performance. In even the simplest game scenarios under normal difficulty, these models frequently resorted to random guessing, with over 95% of key actions requiring hint usage. The inefficiency translated into significantly inflated action sequences and minimal task completion, leading us to our conclusion that 7B is the minimum viable scale for meaningful evaluation in our benchmark.
Model Hints Used Total Steps Early Exit Progress (%)
Qwen-2.5-7B-Instruct 22.6 1329.8 5.93
Qwen-2.5-Coder-7B 22.4 1339.8 12.99
Qwen-2.5-Math-7B 45.4 2384.4 0.00
Table 6: Performance of Domain-Specific Models on EscapeAgent Benchmark
To examine domain specialization effects, we benchmarked two variants from the Qwen-2.5 model series, focused respectively on coding and mathematical tasks. As shown in Table 6, the Qwen-2.5-Coder-7B achieved performance close to the general-purpose 7B-Instruct model, indicating effective domain adaptation without significant loss of reasoning flexibility. In contrast, the Qwen-2.5-Math-7B model struggled, failing to initiate or complete key actions autonomously. This suggests that excessive specialization, such as strict math alignment, may hinder the model’s ability to generalize and engage in creative, multi-step tasks.
6.2 Ablation Study on Key Modules
To evaluate the individual contributions of the Foresight and Reflection modules in EscapeAgent, we conducted an ablation study using two modified agent variants:
  • EscapeAgent (only Foresight): Disables the task list maintained by the Reflection module, allowing action proposals based solely on environmental observation.
  • EscapeAgent (only Reflection): Disables the Foresight module, relying instead on the task list to guide decision-making.
Model Hints Used Total Steps Early Exit Progress (%)
GPT-4o, EscapeAgent 5.00 452.00 44.55
GPT-4o, only Reflection 6.75 570.08 37.00
GPT-4o, only Foresight 7.17 593.92 48.89
GPT-4o, BaseAgent 9.83 707.33 19.85
Llama-3.1-70B, EscapeAgent 7.17 624.67 28.31
Llama-3.1-70B, only Reflection 11.33 878.67 28.78
Llama-3.1-70B, only Foresight 11.00 877.25 22.29
Llama-3.1-70B, BaseAgent 14.67 981.42 17.28
Table 7: Ablation results on GPT-4o and Llama-3.1-70B-Instruct backbones.
We benchmarked these variants alongside the full EscapeAgent and the BaseAgent across all game scenarios with normal difficulty. Across both GPT-4o and Llama-3.1-70B in Table 7, each module independently improves performance relative to the BaseAgent. The full EscapeAgent consistently achieves the best results, confirming that the Foresight and Reflection modules are complementary. Notably, Reflection alone often outperforms Foresight alone, likely because the Foresight module relies on the task list to avoid redundant or failed actions. These results support the conclusion that combining both modules is essential to fully realize EscapeAgent’s potential.